OpenStudy (anonymous):

Choose one of the factors of x15 – 64 x5 – 4 x10 – 4x5 + 16 x10 – 4 x15 + 8x + 16

OpenStudy (cwrw238):

i think its x^5 - 4 - ill try long division

OpenStudy (anonymous):

we can write x^15-64 = (x^5)^3-4^3 now according to the identity a^3-b^3=(a-b)(a^2+ab+b^2) so x^15-64= (x^5-4)(x^10+4x^5+16) therefore x^5-4 is a factor and the answer.

OpenStudy (cwrw238):

yes long division gives (x^5 - 4)(x^10 + 4x^5 + 16)